OpenStudy (anonymous):

graph the line that goes through the point (-4,5) and has the slope 4

OpenStudy (dangerousjesse):

\[\large\frac{ -4-x }{ 5-x }=4\]

OpenStudy (dangerousjesse):

Oops \[\large\frac{-4-x}{5-y}\]